There are two identical sphere A and B . Now Q established on each sphere based on mutual induction . There is third identical neutral sphere C. Now sphere C is first brought in contact with A and separated ,then brought in contact with B and separated after this. What will be charge on C?

Grade:12

1 Answers

Khimraj
5 years ago
initial charge on A and B = Q
charge on C after brought in contact with A = (Q+0)/2 = Q/2.
charge on C after brought in contact with B = (Q + 0.5Q)/2 = 3Q/4